OpenStudy (anonymous):

intigrate ( e^2x) / (e^2x +3)

OpenStudy (sidsiddhartha):

use substitution take\[e^{2x}=u\]

OpenStudy (sidsiddhartha):

then \[2e^{2x}dx=du\] \[e^{2x}dx=du/2\] now just substitute ok??

OpenStudy (sidsiddhartha):

\[1/2\int\limits_{}^{}\frac{ du }{ u+3}\] now can u do it?

OpenStudy (sidsiddhartha):

so what is ur answer?